OpenStudy (anonymous):

What is the midpoint of (–4, 4) and (2, –4)? Answer (–2, 0) (–6, 8) (0, 0) (–1, 0)

OpenStudy (anonymous):

lol i get so confused with these though.. so yeah

OpenStudy (anonymous):

take the average of the x's and the average of the y's in other words add and divide by 2

OpenStudy (anonymous):

\[\frac{-4+2}{2}=\frac{-2}{2}=-1\] and \[\frac{4+(-4)}{2}=\frac{0}{2}=0\] so answer is \[(-1,0)\]

OpenStudy (anonymous):

wow thats easier than i thought haha! thank you(:

OpenStudy (anonymous):

yes it is straightforward. yw